OpenStudy (anonymous):

It takes 30 men to unload 600kg in 3hrs. How many men will it take to unload1600kg in 2hours?

OpenStudy (anonymous):

30 men would unload 200kg in one hour, and one man would unload 200/30=20/3 kg per hour 1600kg in two hours means unloading 800kg per hour. 800kg=2400/3 kg= (120*20)/3 kg Meaning you'd need 120 men